# A Coreset Selection of Coreset Selection Literature: Introduction and Recent Advances

### v1

@@ VI-A Outliers
Random 21.0 $\pm$ 0.3 30.8 $\pm$ 0.6 36.7 $\pm$ 1.7 64.5 $\pm$ 1.1 75.7 $\pm$ 2.0 87.1 $\pm$ 0.5 90.2 $\pm$ 0.3 92.1 $\pm$ 0.1 93.3 $\pm$ 0.2 94.0 $\pm$ 0.2 95.2 $\pm$ 0.1 95.6 $\pm$ 0.1
Herding [37] 20.2 $\pm$ 2.3 27.3 $\pm$ 1.5 34.8 $\pm$ 3.3 51.0 $\pm$ 3.1 63.5 $\pm$ 3.4 74.1 $\pm$ 2.5 80.1 $\pm$ 2.2 85.2 $\pm$ 0.9 88.0 $\pm$ 1.1 89.8 $\pm$ 0.9 94.6 $\pm$ 0.4 95.6 $\pm$ 0.1
k-Center Greedy [22] 18.5 $\pm$ 0.3 26.8 $\pm$ 1.2 31.1 $\pm$ 1.2 51.4 $\pm$ 2.1 75.8 $\pm$ 2.4 87.0 $\pm$ 0.3 90.9 $\pm$ 0.4 92.8 $\pm$ 0.1 93.9 $\pm$ 0.2 94.1 $\pm$ 0.1 95.4 $\pm$ 0.1 95.6 $\pm$ 0.1
Forgetting [47] 21.4 $\pm$ 0.5 29.8 $\pm$ 1.0 35.2 $\pm$ 1.6 52.1 $\pm$ 2.2 67.0 $\pm$ 1.5 86.6 $\pm$ 0.6 91.7 $\pm$ 0.3 93.5 $\pm$ 0.2 94.1 $\pm$ 0.1 94.6 $\pm$ 0.2 95.3 $\pm$ 0.1 95.6 $\pm$ 0.1
GraNd [50] 17.7 $\pm$ 1.0 24.0 $\pm$ 1.1 26.7 $\pm$ 1.3 39.8 $\pm$ 2.3 52.7 $\pm$ 1.9 78.2 $\pm$ 2.9 91.2 $\pm$ 0.7 93.7 $\pm$ 0.3 94.6 $\pm$ 0.1 95.0 $\pm$ 0.2 95.5 $\pm$ 0.2 95.6 $\pm$ 0.1
CAL [60] 22.7 $\pm$ 2.7 33.1 $\pm$ 2.3 37.8 $\pm$ 2.0 60.0 $\pm$ 1.4 71.8 $\pm$ 1.0 80.9 $\pm$ 1.1 86.0 $\pm$ 1.9 87.5 $\pm$ 0.8 89.4 $\pm$ 0.6 91.6 $\pm$ 0.9 94.7 $\pm$ 0.3 95.6 $\pm$ 0.1
DeepFool [27] 17.6 $\pm$ 0.4 22.4 $\pm$ 0.8 27.6 $\pm$ 2.2 42.6 $\pm$ 3.5 60.8 $\pm$ 2.5 83.0 $\pm$ 2.3 90.0 $\pm$ 0.7 93.1 $\pm$ 0.2 94.1 $\pm$ 0.1 94.8 $\pm$ 0.2 95.5 $\pm$ 0.1 95.6 $\pm$ 0.1
Craig [70] 22.5 $\pm$ 1.2 27.0 $\pm$ 0.7 31.7 $\pm$ 1.1 45.2 $\pm$ 2.9 60.2 $\pm$ 4.4 79.6 $\pm$ 3.1 88.4 $\pm$ 0.5 90.8 $\pm$ 1.4 93.3 $\pm$ 0.6 94.2 $\pm$ 0.2 95.5 $\pm$ 0.1 95.6 $\pm$ 0.1
GradMatch [71] 17.4 $\pm$ 1.7 25.6 $\pm$ 2.6 30.8 $\pm$ 1.0 47.2 $\pm$ 0.7 61.5 $\pm$ 2.4 79.9 $\pm$ 2.6 87.4 $\pm$ 2.0 90.4 $\pm$ 1.5 92.9 $\pm$ 0.6 93.2 $\pm$ 1.0 93.7 $\pm$ 0.5 95.6 $\pm$ 0.1
Glister [78] 19.5 $\pm$ 2.1 27.5 $\pm$ 1.4 32.9 $\pm$ 2.4 50.7 $\pm$ 1.5 66.3 $\pm$ 3.5 84.8 $\pm$ 0.9 90.9 $\pm$ 0.3 93.0 $\pm$ 0.2 94.0 $\pm$ 0.3 94.8 $\pm$ 0.2 95.6 $\pm$ 0.2 95.6 $\pm$ 0.1
Facility Location [97] 22.3 $\pm$ 2.0 31.6 $\pm$ 0.6 38.9 $\pm$ 1.4 60.8 $\pm$ 2.5 74.7 $\pm$ 1.3 85.6 $\pm$ 1.9 91.4 $\pm$ 0.4 93.2 $\pm$ 0.3 93.9 $\pm$ 0.2 94.5 $\pm$ 0.3 95.5 $\pm$ 0.2 95.6 $\pm$ 0.1
GraphCut [97] 24.3 $\pm$ 1.5 34.9 $\pm$ 2.3 42.8 $\pm$ 1.3 65.7 $\pm$ 1.2 76.6 $\pm$ 1.5 84.0 $\pm$ 0.5 87.8 $\pm$ 0.4 90.6 $\pm$ 0.3 93.2 $\pm$ 0.3 94.4 $\pm$ 0.3 95.4 $\pm$ 0.1 95.6 $\pm$ 0.1
@@
TABLE III: Coreset selection robustness performances on CIFAR10 with 10% corrupted labels (i.e., 10% of total samples, randomly selected, have been reassigned to new class labels).
@@ VII Applications
Coreset selection has found applications in various domains of machine learning, spanning computer vision [98, 74, 99, 100, 101], natural language processing [102, 103, 104, 105], and quantum machine learning [106, 107, 108]. In the following sections, we examine specific fields where coreset selection has made interesting contributions.
@@ VII-C Dataset Distillation
For instance, He et al. [139] apply coreset selection directly to the distilled dataset by combining a logit-based prediction-error criterion—aiming to retain “easy” samples—with a Rademacher-complexity measure that maintains class balance. Separately, Xu et al. [140] propose a complementary pipeline in which coreset selection, guided by empirical loss, prunes real data before the distillation stage and is further enhanced by a dynamic pruning mechanism during the distillation process.
@@ VII-G Machine Unlearning
Coreset selection can play a crucial role in this process by identifying the most important data points to either retain or remove, thereby minimizing the impact on the model’s overall performance while improving the cost of unlearning [158]. Utility Preserving Coreset Selection (UPCORE) is a framework introduced by [159] for data selection that minimizes the performance degradation during unlearning, focusing on which data points to remove (i.e, the ”forget set”) but also leveraging a coreset selection algorithm to identify which samples to retain.
@@ VII-H Self-supervised Learning
Using coresets in self-supervised learning can improve the efficiency and effectiveness of SSL pre-training. This is critical when training on large open sets (e.g., ImageNet) for domain-specific tasks (e.g., fine-grained classification) where distribution mismatch between the open set and target dataset can degrade performance. Kim et al. [160] propose Simcore, a technique that samples a coreset that minimizes the latent-space distance between the open-set and target dataset. Singh et al. [161] utilize Bloom filters for rapid retrieval of the coreset from the open set and integrate it with Simcore to achieve a reduction in sampling time. RETRIEVE [76] achieves a speedup of around 3$\times$ in the traditional SSL setting.
VII-I Large Language Models
Using coresets in self-supervised learning (SSL) enhances both the efficiency and effectiveness of pre-training, particularly when dealing with large open sets (e.g., ImageNet) for domain-specific tasks such as fine-grained classification. In such scenarios, a distribution mismatch between the open set and the target dataset can negatively impact performance. Kim et al. [160] introduce Simcore, a method that selects a coreset to minimize the latent-space distance between the open-set and target dataset. Similarly, RETRIEVE [76] accelerates traditional SSL training by approximately 3× speedup.
@@ VIII Discussion and Future Directions
On the other hand, training-based approaches, particularly those relying on loss signals, gradient information, and submodular formulations, can yield competitive or even state-of-the-art performance, albeit at a higher computational cost. The recent development in active learning and blind coreset selection further broadens the field, approaching fully unlabeled environments.
@@ VIII-A Discussion
Coverage vs. Difficulty. Methods that prioritize “hard” samples often improve decision boundaries but risk ignoring “easy” or “redundant” data that might be crucial for overall distribution coverage. Future work may consider hybrid approaches that adaptively select from both easy and hard examples, especially in dynamic or incremental learning settings, similar to Zheng et al. [10].
Blind or Label-Free Approaches. Label-free frameworks like ELFS and ZCore offer a promising alternative for scenarios where supervision is limited or nonexistent. However, the quality of their derived coresets hinges on the quality of pre-trained encoders (or pseudo-labeling), which may degrade under domain shifts or for highly specialized tasks.
Influence of Dataset Size and Pruning Ratio. As indicated by recent theoretical and empirical studies, the optimal coreset selection strategy may depend on whether $\mathcal{T}$ is small ($\sim\!1\mathrm{k}$$10\mathrm{k}$ samples) or extremely large ($\sim\!1\mathrm{M}$$10\mathrm{M}$). While large datasets often contain heavy redundancy, smaller ones might demand strategies that retain more “easy” or high-density examples for stable generalization. The same paradigm shift goes for low and high pruning ratios.
Evaluation and Benchmarking. Different works evaluate coreset selection on varying metrics and apply different post-processing (e.g., fine-tuning for 100 vs. 300 epochs). This complicates direct comparisons. Comprehensive benchmarks, possibly spanning multiple architectures and data modalities (image, text, speech), remain an open challenge, while DeepCore [15] offers an initial step in this direction.
@@ VIII-B Future Research Directions
In this section, we want to give a short summary of possible research avenues that can be explored in the future.
Adaptive and Continual Coresets. An important direction is continual learning or streaming scenarios, where new data arrives incrementally. Approaches must update the coreset without full retraining or batch-level re-computation. Combining online gradient matching or submodular updates with incremental embedding updates is a rich research frontier.
Fairness and Robustnes. Beyond mere accuracy, future coreset research can incorporate fairness constraints (ensuring demographic representation) or robust objectives (countering adversarial or noisy labels). Submodular or bilevel methods could be extended to optimize these multicriteria.
Self-Evolving Coresets with Meta-Learning. A highly ambitious direction is to design meta-learning frameworks that continuously refine their subset selection strategy based on the data’s evolving characteristics. Instead of defining a fixed scoring or submodular criterion from the outset, the system would learn how to pick coresets for new tasks or distributions by observing the performance and training dynamics on prior tasks.
Multi-Modal Data. Most coreset research focuses on single-modality data (e.g., images). Extending coverage or gradient-based selection to multi-modal tasks (e.g., image and text) is non-trivial, as each modality’s representation must be jointly considered.
@@ IX Conclusion
This survey has traced the evolution of coreset selection from its geometric foundations to more sophisticated training-oriented and blind methods. By examining core approaches - random and geometry-based sampling, gradient matching, submodular formulations, and bilevel optimization - this work reveals how different strategies strike varied balances between computational efficiency, label requirements, and approximation accuracy.
From a birds-eye view, we highlight how both small and large datasets can be pruned effectively, though the preferred criteria may shift as sample sizes grow or task complexities change. Furthermore, emerging perspectives on coverage, margin-based difficulty, and meta-learning provide broader frameworks for understanding what genuinely constitutes the most “informative” data.

### v2

@@ VI Theoretical Remarks
[…]
Figure 12: Coreset selection performances on CIFAR-10 with 10% corrupted labels, which means that 10% of samples have been reassigned to new class labels [165]. Methods are grouped by our taxonomy and share the same color. Dashed lines distinguish methods within the same category. Most methods show a sharp drop at mid-range pruning ratios (10%–50%). This is most pronounced for geometry-based methods, gradient matching approaches, and GraNd (gradient norm). Forgetting, CAL, and GraphCut remain more stable in comparison to the other methods. Concrete values can be found in Table VI. Methods were evaluated based on DeepCore [43] under identical settings.
A key takeaway is that pruning can break traditional power-law scaling. While random subsampling follows the expected decay $\beta$, a well-pruned dataset can achieve an improved scaling exponent $\beta{\prime}$, leading to significant improvements.
VI-A Implications for Model Design and Loss Construction
Theoretical discussions of pruning often stop at error-vs.-data scaling (e.g., Equation 44), but in practice coreset selection also changes the training signal geometry: it modifies (i) the distribution of margins/difficulties, (ii) the gradient-noise statistics, and (iii) the degree of redundancy. These shifts have concrete implications for model capacity, loss choices, and optimization hyperparameters.
Model design
A useful lens is that pruning changes the dataset’s effective sample size (ESS), which depends on how non-uniform the subset is. If samples in a coreset are used with weights $\{w_{i}\}_{i\in\mathcal{S}}$, a standard proxy for statistical efficiency is
$\mathrm{ESS}(\mathcal{S})\;\triangleq\;\frac{\big(\sum_{i\in\mathcal{S}}w_{i}\big)^{2}}{\sum_{i\in\mathcal{S}}w_{i}^{2}}\;\;\leq\;\;|\mathcal{S}|.$ (46)
Highly skewed selection (e.g., hardest-first scoring without smoothing) can yield $\mathrm{ESS}\ll|\mathcal{S}|$, meaning the model effectively trains on fewer independent samples. This predicts a practical failure mode: large-capacity models overfit more easily under aggressive, high-variance pruning. Quantitatively, if the original training setup was near the interpolation/over-parameterized regime, a conservative guideline is to keep the ratio $\frac{P}{\mathrm{ESS}}$ (parameters per effective sample) from increasing drastically when moving from $\mathcal{T}$ to $\mathcal{S}$. When $\mathrm{ESS}$ collapses, one may need stronger regularization (weight decay, augmentation, dropout) or smaller models to maintain generalization.
Loss Construction
Coreset objectives implicitly reshape the distribution of margins and label noise: hard-sample selectors (GraNd/EL2N, boundary-based) bias toward low-margin, high-loss points, while coverage/submodular selectors bias toward distributional support. This matters because standard cross-entropy is sensitive to mislabeled outliers. Hence, under noisy or outlier-heavy pruning, robustification becomes more important. Two practical, quantitative knobs are: (i) label smoothing (e.g., replace one-hot $y$ with $(1-\epsilon)y+\epsilon/K$), and (ii) loss clipping / robust losses that reduce the influence of extreme-loss samples.
VI-B Concerning Other Modalities
While coreset selection is applicable across modalities (see Section VII), most comparable empirical evidence is concentrated in image classification benchmarks, where selection pipelines, architectures, and evaluation protocols are relatively standardized (e.g., fixed training recipes, shared backbones, and consistent metrics), enabling controlled cross-method comparisons. In contrast, large-scale vision (e.g., ImageNet-scale) and NLP pretraining introduce confounds (tokenization and sequence length effects, heavy deduplication/filtering pipelines, domain-mixing, and substantial training-budget sensitivity) that make “apples-to-apples” benchmarking across coreset objectives considerably harder and often entangled with data quality filtering rather than subset selection alone.
VI-C Outliers
Another interesting aspect of coreset selection, often overlooked, is the role of outliers [120, 6]. Outliers can either provide valuable decision boundary information or introduce noise and perturbations during the training. The prior handling of outliers, e.g., with deep SVDD [120, 177] or contrastive score-based OOD [142], can significantly improve the performance of any coreset selection method [147].
A practical caveat is that, in many real-world settings, the presence (and rate) of outliers or label noise is unknown a priori. Most coreset selection methods are developed and tuned under an implicit clean-data assumption. When this assumption is violated, selection scores can become dominated by mislabeled or anomalous samples, leading to highly skewed subsets and drastic performance drops. This motivates treating outlier/OOD filtering not as an optional pre-processing step, but as a robustness-critical component of any future coreset pipeline, especially under aggressive pruning.
We provide in Figure 12 and Table VI additional robustness experiments to highlight the importance of outliers. In this experiment, we corrupted 10% of the labels, i.e., reassigning 10% of samples to new class labels [165]. The selection of corrupted labels was repeated during each coreset sampling procedure. To make the empirical comparisons interpretable, we emphasize that the results follow an identical training/evaluation protocol across methods. Concretely, we evaluated a ResNet-18 with SGD (initial learning rate $0.1$ with cosine decay, momentum $0.9$, weight decay $5\times 10^{-4}$), trained for $200$ epochs with standard CIFAR-10 augmentation (random crop with $4$-pixel padding and random horizontal flip). Training-oriented selectors compute their scores from a partially trained network (a $10$-epoch pretraining stage for scoring), and reported numbers are averaged over $5$ random seeds (mean $\pm$ std), as provided by the benchmark.
As a result, one can see that most sampling methods get negatively affected (red), especially geometry-based, gradient-matching-based methods, and GraNd (gradient norm). Interestingly, forgetting, CAL, and GraphCut remain more stable in comparison to the other methods. Furthermore, CAL confirms Equation 44, as its $\alpha=10\%$ pruned dataset outperforms the training on the full (corrupted) dataset. Also, forgetting and GraphCut were able to achieve better results for some cases.
As this experiment highlights, future work could further explore automated ways prior to coreset selection to filter outliers dynamically. For instance, Zheng et al. [173] leveraged the area under margin [113], which captures how far a sample is from the decision boundary over multiple training iterations. Other innovations might include classical outlier detection principles, as shown in HyperCore [101].
@@ VI-C Outliers
TABLE VI: Robustness on CIFAR10 with 10% corrupted labels (i.e., 10% of total samples, randomly selected, have been reassigned to new class labels). Methods were evaluated based on DeepCore [43] under identical settings.
@@ VII Applications
Coreset selection has found applications in various domains, spanning computer vision [79, 176, 90, 58, 77], natural language processing [7, 59, 168, 140], and quantum machine learning [114, 53, 156]. In the following sections, we examine where coreset selection has made interesting contributions.
Across applications, the key difference is which signal is available for selection: training-free pipelines operate in a fixed embedding space (often from foundation models) to optimize coverage/diversity, training-oriented methods leverage partial training dynamics (loss/gradients/margins) for model-specific utility, and label-free variants replace ground-truth labels with representation proxies or pseudo-labels (scalable but brittle under domain shift).
@@ VII-C Dataset Distillation
For instance, He et al. [49] apply coreset selection directly to the distilled dataset by combining a logit-based prediction-error criterion (aiming to retain “easy” samples) with a Rademacher-complexity measure that maintains class balance. Separately, Xu et al. [155] propose a complementary pipeline in which coreset selection, guided by empirical loss, prunes real data before the distillation stage and is further enhanced by a dynamic pruning mechanism during the distillation process.
@@ VII-G Machine Unlearning
Coreset selection can play a crucial role in this process by identifying the most important data points to either retain or remove, thereby minimizing the impact on the model’s overall performance while improving the cost of unlearning [16]. Utility Preserving Coreset Selection (UPCORE) is a framework introduced by [109] for data selection that minimizes the performance degradation during unlearning, focusing on which data points to remove (i.e., the ”forget set”) but also leveraging a coreset selection algorithm to identify which samples to retain.
@@ VII-H Self-supervised Learning
[…]
Another emerging direction is to increasingly use foundation-model embedding spaces, often trained self-supervised, as the default geometry for training-free and label-free selection, e.g., DINO [19], or DINOv2 [108] representations, which tend to be more semantically aligned than task-trained backbones.
VII-I LLM Pretraining
LLM pretraining corpora span trillions of tokens, making coreset selection a practical necessity: RedPajama [148] approximates the LLaMA pipeline by removing low-quality Common Crawl documents via score-based signals (e.g., CCNet perplexity sorting to drop the high-perplexity tail), while Dolma [128] curates roughly 200TB of raw text into a cleaned 11TB (about 3T tokens) dataset through label-free quality filters (e.g., boilerplate stripping and removal of nonsensical/offensive text). Beyond quality filtering, training-free geometry-based selection is central for redundancy reduction at web scale: SlimPajama [48] applies global MinHash-LSH soft deduplication to remove near-duplicate passages, eliminating about 49.6% of 1.2T tokens and yielding a 627B-token subset with higher information density, while corpus mixing across modalities/sources (code, books, web, scientific text) in RedPajama and Dolma can be interpreted as coarse cluster coverage for diversity. Training-based scoring further aligns selection with downstream learning signals: RefinedWeb [111] demonstrates that aggressive filtering and deduplication of Common Crawl using learned criteria (including “Gopher”-style rules to discard gibberish and overly simplistic content) produces a high-quality web corpus (with a released 600B-token extract) that improves model outcomes, and large-scale dataset construction efforts such as The Pile [39] and ROOTS [78] emphasize broad domain coverage (implicitly echoing submodular/coverage objectives) even when explicit submodular optimization is not directly applied at trillion-token scales.
@@ VIII Discussion and Future Directions
On the other hand, training-based approaches, particularly those relying on loss signals, gradient information, and submodular formulations, can yield competitive or even state-of-the-art performance, albeit at a higher computational cost. The recent development in active learning and label-free coreset selection further broadens the field, approaching fully label-free environments.
@@ VIII-A Discussion
Coverage vs. Difficulty. A recurring practical question is: How do we keep decision-boundary information and preserve distributional coverage under aggressive pruning? Hard-sample scoring (GraNd/EL2N, boundary-based) tends to sharpen boundaries but can collapse coverage and over-select outliers, especially at high pruning ratios. Conversely, coverage-centric families (submodular facility-location/graph-cut, $k$-Center) preserve modes but may miss boundary refinement. This suggests hybrids that explicitly mix coverage (training-free/submodular) with difficulty (training-oriented scores), potentially with a budget split that adapts with dataset size and pruning ratio, similar to Zheng et al. [173].
Label-Free Approaches. In many real settings, the question is: Can we select a useful subset without labels while remaining robust to distribution shift? Purely label-free pipelines like ELFS and ZCore (clustering/pseudo-labeling or VLM-based embeddings + coverage) scale well, but their failure mode is brittle representations: pseudo-label noise and embedding misalignment can dominate the subset. In this regime, coverage/submodular selection in transferable embeddings is typically safer than pseudo-label-driven “hardness” scoring, and robustness can be improved by coupling selection with simple OOD/outlier filters or agreement signals (e.g., multi-view consistency).
Influence of Dataset Size and Pruning Ratio.
A practical selection question is: Which family should we use when $|\mathcal{T}|$ and $(1-\alpha)$ change? Empirically, at extreme compression (very small $(1-\alpha)$), coverage-oriented selectors and submodular objectives are often preferable to avoid mode dropping, whereas at moderate compression, training-oriented scoring can better target model-specific utility. This also interacts with noise: hardest-first criteria can become outlier-dominated when labels are corrupted, suggesting either robustified scoring (smoothing/clipping) or coverage-first selection when noise is suspected.
Evaluation and Benchmarking. A key open question is: Which coreset families transfer across architectures and remain reliable when the downstream model or domain changes? Training-oriented selectors can be tightly coupled to the scoring model and training recipe, while embedding-based training-free/submodular selectors may transfer better if the representation is strong. This calls for benchmarks that explicitly vary (i) architecture, (ii) fine-tuning protocol, and (iii) domain shift/noise, and report both selection cost and downstream utility rather than accuracy alone.
@@ VIII-B Future Research Directions
We summarize future directions as concrete research questions and typical deployment scenarios, and indicate which method families appear most suitable in each case.
Adaptive and Continual Coresets. How can we maintain a small replay buffer that remains representative as the stream distribution drifts? This scenario favors online coverage (incremental $k$-Center/submodular updates in an evolving embedding space) combined with lightweight training signals (periodic scoring refresh, forgetting-style dynamics) to prevent replay collapse onto stale modes. Bilevel objectives are attractive when a small validation buffer exists per phase, but their cost motivates approximate, amortized updates.
Fairness and Robustness. Can we design selectors that avoid outlier domination and remain stable under corrupted labels or adversarial contamination? Submodular coverage/diversity is a strong baseline because it spreads budget across modes, while bilevel selection can explicitly target robust validation objectives when trustworthy validation data exists. A promising direction is to integrate robust scoring (clipped losses, uncertainty-aware weighting) with explicit OOD filtering prior to selection, and to evaluate selectors by worst-group or worst-slice performance, not only mean accuracy.
Self-Evolving Coresets with Meta-Learning. Can we meta-learn selection policies that generalize across datasets, model families, and training recipes? This naturally aligns with (i) amortized scoring models that predict utility from representations, and (ii) bilevel/meta-learning formulations where the outer objective measures transfer to new tasks. A key challenge is avoiding overfitting to a single benchmark: future work should evaluate cross-architecture transfer and robustness to changes in optimizer, augmentation, and label quality.
Multi-Modality. For VLM/LLM fine-tuning, what subset best preserves coverage over domains, styles, and rare capabilities while controlling redundancy? Here, training-free selection in joint embedding spaces (CLIP-like) and submodular objectives over multimodal similarity are natural first choices, because they directly target diversity/coverage without requiring task labels. When downstream preferences are known (e.g., instruction-following, safety), bilevel/validation-driven selection becomes compelling, using held-out task metrics as the outer objective.
@@ IX Conclusion
This survey has traced the evolution of coreset selection from its geometric foundations to more sophisticated training-oriented and label-free methods. By examining core approaches - random and geometry-based sampling, gradient matching, submodular formulations, and bilevel optimization - this work reveals how different strategies strike varied balances between computational efficiency, label requirements, and accuracy.
From a bird’s-eye view, we highlight how both small and large datasets can be pruned effectively, though the preferred criteria may shift as sample sizes grow or task complexities change. Furthermore, emerging perspectives on coverage, margin-based difficulty, and meta-learning provide broader frameworks for understanding what genuinely constitutes the most “informative” data.